# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v1

@@ Abstract
Classical reactive synthesis approaches aim to synthesize a reactive system that always satisfies a given specifications. These approaches often reduce to playing a two-player zero-sum game where the goal is to synthesize a winning strategy. However, in many pragmatic domains, such as robotics, a winning strategy does not always exist, yet it is desirable for the system to make an effort to satisfy its requirements instead of “giving up”. To this end, this paper investigates the notion of admissible strategies, which formalize “doing-your-best”, in quantitative reachability games. We show that, unlike the qualitative case, quantitative admissible strategies are history-dependent even for finite payoff functions, making synthesis a challenging task. In addition, we prove that admissible strategies always exist but may produce undesirable optimistic behaviors. To mitigate this, we propose admissible winning strategies, which enforce the best possible outcome while being admissible. We show that both strategies always exist but are not memoryless. We provide necessary and sufficient conditions for the existence of both strategies and propose synthesis algorithms. Finally, we illustrate the strategies on gridworld and robot manipulator domains.
@@ 1 Introduction
Reactive Synthesis is the problem of automatically generating reactive systems from logical specifications, first proposed by Church (1963). Its applications span a wide range of domains, including robotics (McMahon et al. 2023; He et al. 2017; Kress-Gazit, Lahijanian, and Raman 2018), program synthesis (Pnueli and Rosner 1989), distributed systems (Filippidis and Murray 2016), formal verification (Kupferman and Vardi 2001), and security (Zhou and Foley 2003). Existing approaches to reactive synthesis usually boil down to computing a strategy over a game between a System (Sys) and an Environment (Env) player, with the goal of finding a winning strategy, which guarantees the Sys player achieves its objectives regardless of the Env player’s moves. In quantitative settings, the game incorporates a payoff requirement as part of the Sys player’s objectives. In many scenarios, however, a winning strategy may not exist, resulting in a failure in the synthesis algorithm. Prior works relax this requirement using the notions of best effort (Aminof et al. 2021) and admissibility (Berwanger 2007; Brandenburger, Friedenberg, and Keisler 2008). Best effort is specific to qualitative games, while admissibility is recently studied in quantitative games under strong assumptions on the Env player, namely, rationality and known objectives (Brenguier et al. 2016). This paper aims to study admissible strategies without these assumptions in quantitative reachability games with a particular focus on applications in robotics.
@@
Consider the example in Fig. 1 with a robot (Sys player) and a human (Env player) operating in shared workspace. The robot is tasked with grasping the bin. Since the human can intervene by moving the bin before the robot completes its grasp, there is no winning strategy that enforces completion of the task under the worst-case Env strategy. However, in such cases it is still desirable for the robot to make an effort to satisfy its requirements instead of “giving up”.
@@
To this end, this paper studies admissible strategies (Faella 2009; Berwanger 2007) in quantitative reachability games without assumptions on rationality and the objectives of the Env player. These games can model quantitative reactive synthesis for finite-behaviors expressed in, e.g., syntactically co-safe Linear Temporal Logic (ltl) (Kupferman and Vardi 2001) and ltl over finite behaviors ($\textsc{ltl}_{f}$) (De Giacomo and Vardi 2013). We show that admissible strategies relax requirement of winning strategies, and always exist. We prove that, unlike the qualitative setting, quantitative admissible strategies are history-dependent even for finite payoff functions. Then, we show that such strategies can produce overly optimistic behaviors, which may be undesirable for robotics applications. To mitigate this, we propose admissible winning strategies, which have the desirable property of enforcing specification satisfaction when possible while being admissible. We prove that, similar to admissible strategies, admissible winning strategies always exist and require finite memory. Then, we provide necessary and sufficient conditions for the existence of both strategies and propose synthesis algorithms. Finally, we provide various robotic examples to show that admissible and admissible winning strategies provide desirable and flexible behaviors without a-priori knowledge of the objectives of other agents.
@@
Our contributions are (i) analysis of admissible strategies in quantitative reachability games without assumptions on rationality and objective of the Env player, including proofs that such strategies always exist, require finite memory, and necessary and sufficient conditions for their existence, (ii) introduction and analysis of novel notion of admissible winning strategies to mitigate over-optimism in admissible strategies, (iii) synthesis algorithms for both strategies, and (iv) illustrative examples on gridworld and manipulation domains, showing emergent behaviors.
Related Work
@@
Several works explore alternatives to winning strategies. Specifically, Faella (2009) investigates various concepts in qualitative games with reachability objectives within a zero-sum framework. They focuse solely on the Sys player’s objective without making any assumptions about the Env player. They use the notion of admissibility to define the notion of best-effort (BE). Aminof et al. (2021); De Giacomo, Parretti, and Zhu (2023) further examine the complexity of synthesizing BE strategies, showing that it can be reduced to standard algorithms, with memoryless strategies being sufficient. In contrast, our work considers quantitative reachability games where the objective is to reach a goal state with minimal total cost. We show that memoryless strategies are insufficient in our context, and our synthesis approach does not reduce to standard algorithms.
@@
The notion of admissiblity has also been explored in normal form games (Brandenburger, Friedenberg, and Keisler 2008; Apt 2011). In qualitative games with logical specifications (extensive form), admissiblity has been investigated for n-player infinite games, where each player has their own objective and is assumed to play admissibly with respective to that objective – referred to as assume admissible (AA). Berwanger (2007) was the first to formalize this notion of AA for games played on graphs. Subsequently, Brenguier, Raskin, and Sassolas (2014); Brenguier, Raskin, and Sankur (2015) establish the complexity and give algorithms for $\omega$-regular objectives. In our settings, we consider reachability games that terminate in finite time. Notably, we make no assumptions about the Env player, i.e, we neither know Env player’s objective nor require them to play admissibly.
@@
The work closest to ours is by Brenguier et al. (2016), who study admissibility in quantitative settings. They extend prior works (Brenguier, Raskin, and Sassolas 2014; Brenguier, Raskin, and Sankur 2015) on infinite duration qualitative games to quantitative objectives. They analyze and give necessary and sufficient conditions for admissible strategies to exists. Unlike their work, in our settings, we consider finite duration games and appropriately define our payoff over finite traces. We show our game is always determined and thus optimal worst-case and cooperative strategies always exist. While our analysis shares some conceptual similarities with theirs, addressing the finite play setting requires a distinct theoretical approach compared to infinite plays. Brenguier et al. (2016) give a sketch of their algorithm based on parity games. We, however, present a detailed yet simpler synthesis algorithm. We also analyze emergent behavior under admissible strategies in robotics settings. We observe that these strategies can be overly optimistic. To address this, we identify the underlying cause and propose the concept of admissible winning strategies to mitigate such optimism.
@@ 2 Problem Formulation
The overarching goal of this work is quantitative reactive synthesis for $\textsc{ltl}_{f}$ or cosafe ltl specifications where satisfaction cannot necessarily be guaranteed. This problem reduces to reachability analysis in quantitative games played between the Sys and Env players (Baier and Katoen 2008). For the sake of generality, we focus on these games.
@@
  • $C:V\times(A_{s}\cup A_{e})\to\mathbb{N}\cup\{0\}$ is the cost (energy) function such that, for every $(v,a)\in V_{s}\times A_{s}$, $C(v,a)>0$, otherwise 00, and
@@
We assume that the game is non-blocking. There is at least one outgoing transition from every state, i.e., $\forall v\in V$ and $\forall a\in A$, $\delta(v,a)\neq\emptyset$. Note that the Env action cost is zero since we are solely interested in the Sys player objectives and action costs, and we make no assumption about the objective of the Env player. Finally, we assume that our transition function is deterministic and injective, i.e., $\delta(v,a)=\delta(v,a^{\prime})\iff a=a^{\prime}$.
The evolution of game $\mathcal{G}$ starts from $v_{0}$ and is played in turns between the Sys and Env player. At state $v\in V_{i}$, where $i\in\{s,e\}$, Player $i$ picks an action $a\in A_{i}$ and incurs cost $C(v,a)$. Then, the game evolves to the next state according to the transition function $\delta(v,a)\in V_{j}$, where $j\neq i$. Then, Player $j$ picks an action, and the process repeats. The game terminates if a goal state in $V_{f}$ is reached. The players choose actions according to a strategy.
@@
Given strategies $\sigma$ and $\tau$, a unique sequence of states, called play and denoted by $P^{v_{0}}(\sigma,\tau)$, is induced from $v_{0}$. A play can be finite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots v_{n}\in V^{*}$ or infinite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots\in V^{\omega}$. We denote by $\operatorname{Plays}^{v}:=\{P^{v}(\sigma,\tau)\mid\sigma\in\Sigma,\;\tau\in\mathrm{T}\}$ the set of plays starting from $v$ under every Sys and Env strategy. $\operatorname{Plays}^{v}(\sigma)$ is the set of plays induced by a fixed strategy $\sigma$ and every Env strategy. We note that a finite play occurs iff a goal state is reached.
@@
Given strategies $\sigma\in\Sigma$ and $\tau\in\mathrm{T}$, total-payoff is defined as the sum of all the action costs given by $C$ along the induced play $P^{v_{0}}(\sigma,\tau)=v_{0}v_{1}\ldots v_{n}$ where $n\in\mathbb{N}\cup\{\infty\}$, i.e.,
$\operatorname{Val}(P^{v_{0}}(\sigma,\tau)):=\sum_{i=0}^{n-1}C(v_{i},a_{i},v_{i+1}),$ (1)
where $a_{i}=\sigma(v_{0}\ldots v_{i})$ if $v_{i}\in V_{s}$, or else $a_{i}=\tau(v_{0}\ldots v_{i})$.
@@
Given Sys strategy $\sigma$ and history $h$, let $h\cdot P^{h}(\sigma,\tau)$ denote a play that extends history $h$, where $\sigma$ and $\tau$ are strategies compatible with $h$. The cooperative value $\operatorname{cVal}(h,\sigma)$ is the payoff of the play such that the Env player plays minimally, i.e.,
$\operatorname{cVal}(h,\sigma)=\min_{\tau\in\mathrm{T}}\,\operatorname{Val}(h\cdot P^{h}(\sigma,\tau)).$ (2)
@@
$\operatorname{aVal}(h,\sigma)=\max_{\tau\in\mathrm{T}}\,\operatorname{Val}(h\cdot P^{h}(\sigma,\tau)).$ (3)
@@
$\operatorname{cVal}(h)=\min_{\sigma\in\Sigma}\operatorname{cVal}(h,\sigma)\;\>\text{and}\;\>\operatorname{aVal}(h)=\min_{\sigma\in\Sigma}\operatorname{aVal}(h,\sigma).$
We say a Sys strategy $\sigma_{win}\in\Sigma$ is winning if $\operatorname{aVal}(h,\sigma_{win})<\infty$. If $\sigma_{win}$ exists, the Sys player can force a visit to $V_{f}$ for all Env strategies, guaranteeing its reachability objective. In classical quantitative reactive synthesis, the interest is in a $\sigma_{win}$ that achieves the optimal $\operatorname{aVal}(v_{0})$.
@@
In many applications, when a winning strategy does not exist, it is desirable for the Sys player to adopt a strategy that ensures the best possible outcome. For game $\mathcal{G}$ in Fig. 1, $\sigma_{win}$ does not exist but it is better for the Sys player to keep trying, i.e., move to $v_{1}$ or $v_{3}$ rather than giving up. To capture this intuition, we turn to the classical notion of dominance.
@@
Definition 5 (Dominance (Brenguier et al. 2016; Berwanger 2007)).
Given two Sys strategies $\sigma,\sigma^{\prime}\in\Sigma$, we say
@@
Brandenburger, Friedenberg, and Keisler (2008) rationalize a player playing admissibly to be doing their best. Thus, the first problem we consider is the synthesis of admissible strategies for Sys player in $\mathcal{G}$.
Problem 1 (Admissible Strategy Synthesis).
Given a 2-player quantitative reachability game $\mathcal{G}$ and energy budget $\mathcal{B}\in\mathbb{N}$, synthesize the set of all admissible strategies $\Sigma_{adm}$ such that, for every $\sigma\in\Sigma_{adm}$, there exists $\tau\in\mathrm{T}$ under which $\operatorname{Val}(P^{v_{0}}(\sigma,\tau))\leq\mathcal{B}$.
Intuitively, $\sigma\in\Sigma_{adm}$ allows the Sys player to do its best without any assumptions about the Env player. In qualitative settings, winning strategies are always admissible (Brenguier, Raskin, and Sankur 2015, Lemma 8). In quantitative settings, however, winning strategies are not necessarily admissible (e.g., $\sigma_{1}$ in the game in Fig. 2). Thus, to compute enforceable admissible strategies, we introduce the notion of admissible winning strategy.
@@
Given budget $\mathcal{B}\in\mathbb{N}$, strategy $\sigma$ is called admissible winning for the Sys player if it is admissible and $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$ if $\operatorname{aVal}(h)<\infty$ then $\operatorname{Val}(h\cdot P^{h}(\sigma,\tau))\leq\mathcal{B}$ for all $\tau\in\mathrm{T}$.
For the game in Fig. 2, $\sigma_{3}$ is admissible but not enforcing as it can not ensure always reaching the goal state. $\sigma_{1}$ is winning but not admissible. $\sigma_{2}$ is admissible winning, which is more desirable than the other two.
Problem 2 (Admissible Winning Strategy Synthesis).
Given a 2-player quantitative reachability game $\mathcal{G}$ and energy budget $\mathcal{B}\in\mathbb{N}$, synthesize the set of all admissible winning strategies $\Sigma^{win}_{adm}$.
In Section 3, we show how to solve Problem 1 by providing necessary and sufficient conditions for a strategy to be admissible. In Section 4, we identify the class of admissible strategies that are admissible winning and give an algorithm to solve Problem 2.

### v2

@@ Abstract
Classical reactive synthesis approaches aim to synthesize a reactive system that always satisfies a given specification. These approaches often reduce to playing a two-player zero-sum game where the goal is to synthesize a winning strategy. However, in many pragmatic domains, such as robotics, a winning strategy does not always exist, yet it is desirable for the system to make an effort to satisfy its requirements instead of “giving up.” To this end, this paper investigates the notion of admissible strategies, which formalize “doing-your-best”, in quantitative reachability games. We show that, unlike the qualitative case, memoryless strategies are not sufficient to capture all admissible strategies, making synthesis a challenging task. In addition, we prove that admissible strategies always exist but may produce undesirable optimistic behaviors. To mitigate this, we propose admissible winning strategies, which enforce the best possible outcome while being admissible. We show that both strategies always exist but are not memoryless. We provide necessary and sufficient conditions for the existence of both strategies and propose synthesis algorithms. Finally, we illustrate the strategies on gridworld and robot manipulator domains.
@@ 1 Introduction
Reactive Synthesis is the problem of automatically generating reactive systems from logical specifications, first proposed by Church [1963]. Its applications span a wide range of domains, including robotics [McMahon et al., 2023; He et al., 2017; Kress-Gazit et al., 2018], program synthesis [Pnueli and Rosner, 1989], distributed systems [Filippidis and Murray, 2016], formal verification [Kupferman and Vardi, 2001], and security [Zhou and Foley, 2003]. Existing approaches to reactive synthesis usually boil down to computing a strategy over a game between a System (Sys) and an Environment (Env) player, with the goal of finding a winning strategy, which guarantees the Sys player achieves its objectives regardless of the Env player’s moves. In quantitative settings, the game incorporates a payoff requirement as part of the Sys player’s objectives. In many scenarios, however, a winning strategy may not exist, resulting in a failure in the synthesis algorithm. Prior works relax this requirement using the notions of best effort [Aminof et al., 2021] and admissibility [Berwanger, 2007; Brandenburger et al., 2008]. Best effort is specific to qualitative games, while admissibility is recently studied in quantitative games under strong assumptions on the Env player, namely, rationality and known objectives [Brenguier et al., 2016]. This paper aims to study admissible strategies without these assumptions in quantitative reachability games with a particular focus on applications in robotics.
@@
Consider the example in Fig. 1 with a robot (Sys player) and a human (Env player) operating in shared workspace. The robot is tasked with grasping the bin. Since the human can intervene by moving the bin before the robot completes its grasp, there is no winning strategy that enforces completion of the task under the worst-case Env strategy. However, in such cases it is still desirable for the robot to make an effort to satisfy its requirements instead of “giving up.”
@@
To this end, this paper studies admissible strategies [Faella, 2009; Berwanger, 2007] in quantitative reachability games without assumptions on rationality and the objectives of the Env player. These games can model quantitative reactive synthesis for finite-behaviors expressed in, e.g., syntactically co-safe Linear Temporal Logic (ltl) [Kupferman and Vardi, 2001] and ltl over finite behaviors ($\textsc{ltl}_{f}$) [De Giacomo and Vardi, 2013]. We show that admissible strategies relax requirement of winning strategies, and always exist. We prove that, unlike the qualitative setting, quantitative admissible strategies are generally history-dependent even for finite payoff functions. Then, we show that such strategies can produce overly optimistic behaviors, which may be undesirable for robotics applications. To mitigate this, we propose admissible winning strategies, which have the desirable property of enforcing specification satisfaction when possible while being admissible. We prove that similar to admissible strategies, admissible winning strategies always exist and may require finite memory. Then, we provide necessary and sufficient conditions for both strategies and propose synthesis algorithms. Finally, we provide various robotic examples to show that admissible and admissible winning strategies provide desirable and flexible behaviors without a-priori knowledge of the objectives of other agents.
@@
Our contributions are fourfold: (i) analysis of admissible strategies in quantitative reachability games without assumptions on rationality and objective of the Env player, including proofs of their existence, finite memory requirement, and necessary and sufficient conditions, (ii) introduction and analysis of novel notion of admissible winning strategies to mitigate over-optimism in admissible strategies, (iii) synthesis algorithms for both strategies, and (iv) illustrative examples on gridworld and manipulation domains, showing emergent behaviors under these strategies.
Related Work.
@@
Several works explore alternatives to winning strategies. Specifically, Faella [2009] investigates various concepts in qualitative games with reachability objectives within a zero-sum framework. They focus solely on the Sys player’s objective without making any assumptions about the Env player. They use admissibility to define the notion of best-effort (BE). Aminof et al. [2021]; De Giacomo et al. [2023] further examine the complexity of synthesizing BE strategies, showing that it can be reduced to standard algorithms, with memoryless strategies being sufficient. In contrast, our work considers quantitative reachability games where the objective is to reach a goal state with minimal total cost. We show that memoryless strategies are insufficient in our context, and our synthesis approach does not reduce to standard algorithms.
@@
The notion of admissiblity has also been explored in normal form games [Brandenburger et al., 2008; Apt, 2011]. In qualitative games with logical specifications (extensive form), admissiblity has been investigated for n-player infinite games, where each player has their own objective and is assumed to play admissibly with respective to that objective – referred to as assume admissible (AA). Berwanger [2007] was the first to formalize this notion of AA for games played on graphs. Subsequently, Brenguier et al. [2014, 2015] establish the complexity and give algorithms for $\omega$-regular objectives. In our settings, we consider reachability games that terminate in finite time. Notably, we make no assumptions about the Env player, i.e., we neither know Env player’s objective nor require them to play admissibly.
@@
The work closest to ours is by Brenguier et al. [2016], who study admissibility in quantitative settings. They extend prior work [Brenguier et al., 2014, 2015] on infinite duration qualitative games to quantitative objectives. They give necessary and sufficient conditions for admissible strategies. Unlike their work, in our setting, we consider finite duration games and appropriately define our payoff over finite traces. We show our game is always determined and thus optimal worst-case and cooperative strategies always exist. While our analysis shares some conceptual similarities with theirs, addressing the finite play setting requires a distinct theoretical approach compared to infinite plays. Brenguier et al. [2016] give a sketch of their algorithm based on parity games. We, however, present a detailed yet simpler synthesis algorithm. We also analyze emergent behavior under admissible strategies in robotics settings. We observe that these strategies can be overly optimistic. To address this, we identify the underlying cause and propose the concept of admissible winning strategies to mitigate such optimism.
@@ 2 Problem Formulation
The overarching goal of this work is quantitative reactive synthesis for $\textsc{ltl}_{f}$ or cosafe ltl specifications where satisfaction cannot necessarily be guaranteed. This problem reduces to reachability analysis in quantitative games played between the Sys and Env players [Baier and Katoen, 2008]. For the sake of generality, we focus on these games.
@@
  • $C:V\times(A_{s}\cup A_{e})\to\mathbb{N}^{0}$ is the cost (energy) function such that, for every $(v,a)\in V_{s}\times A_{s}$, $C(v,a)>0$, otherwise 00, and
@@
We assume that the game is non-blocking. There is at least one outgoing transition from every state, i.e., $\forall v\in V,\exists a\in A$ s.t. $\delta(v,a)\neq\emptyset$. Note that the Env action cost is zero since we are solely interested in the Sys player objectives (action costs) and make no assumption about the objective of the Env player. Finally, we assume that our transition function is deterministic and injective, i.e., $\delta(v,a)=\delta(v,a^{\prime})$ iff $a=a^{\prime}$.
The evolution of game $\mathcal{G}$ starts from $v_{0}$ and is played in turns between the Sys and Env player. At state $v\in V_{i}$, where $i\in\{s,e\}$, Player $i$ picks an action $a\in A_{i}$ and incurs cost $C(v,a)$. Then, the game evolves to the next state according to the transition function $\delta(v,a)\in V_{j}$, where $j\neq i$. Then, Player $j$ picks an action, and the process repeats. The game terminates if a goal state in $V_{f}$ is reached. For the remainder of the paper, all definitions are provided with respect to the game $\mathcal{G}$. For brevity, we omit explicitly restating this context.
The players choose actions according to a strategy.
@@
Given strategies $\sigma$ and $\tau$, a unique sequence of states, called play and denoted by $P^{v_{0}}(\sigma,\tau)$, is induced from $v_{0}$. A play can be finite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots v_{n}\in V^{*}$ or infinite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots\in V^{\omega}$.
We denote by $\operatorname{Plays}^{v}:=\{P^{v}(\sigma,\tau)\mid\sigma\in\Sigma,\;\tau\in\mathrm{T}\}$ the set of plays starting from $v$ under every Sys and Env strategy. $\operatorname{Plays}^{v}(\sigma)$ is the set of plays induced by a fixed strategy $\sigma$ and every Env strategy. We note that a finite play occurs iff a goal state is reached.
@@
Given strategies $\sigma\in\Sigma$ and $\tau\in\mathrm{T}$, total payoff is defined as the sum of all the action costs given by $C$ along the induced play $P^{v_{0}}(\sigma,\tau)=v_{0}v_{1}\ldots v_{n}$ where $n\in\mathbb{N}\cup\{\infty\}$, i.e.,
$\operatorname{Val}(P^{v_{0}}(\sigma,\tau)):=\sum_{i=0}^{n-1}C(v_{i},a_{i}),$ (1)
where $a_{i}=\sigma(v_{0}\ldots v_{i})$ if $v_{i}\in V_{s}$, else $a_{i}=\tau(v_{0}\ldots v_{i})$.
@@
Given $h$, Sys ($\sigma$) and Env ($\tau$) strategies compatible with $h$, let $P^{h}(\sigma,\tau)$ denote a play that extends history $h$ The cooperative value $\operatorname{cVal}(h,\sigma)$ is the payoff of the play such that the Env player plays minimally, i.e.,
$\operatorname{cVal}(h,\sigma)=\inf_{\tau\in\mathrm{T}}\,\operatorname{Val}(P^{h}(\sigma,\tau)).$ (2)
@@
$\operatorname{aVal}(h,\sigma)=\sup_{\tau\in\mathrm{T}}\,\operatorname{Val}(P^{h}(\sigma,\tau)).$ (3)
@@
$\operatorname{cVal}(h)=\inf_{\sigma\in\Sigma}\operatorname{cVal}(h,\sigma)\;\>\text{and}\;\>\operatorname{aVal}(h)=\inf_{\sigma\in\Sigma}\operatorname{aVal}(h,\sigma).$
For a given $h$, we say Sys strategy $\sigma_{win}\in\Sigma$ is winning if $\sigma_{win}$ is compatible with $h$ and $\operatorname{aVal}(h,\sigma_{win})<\infty$. If $\sigma_{win}$ exists, the Sys player can force a visit to $V_{f}$ for all Env strategies, guaranteeing its reachability objective. In classical quantitative reactive synthesis, the interest is in a $\sigma_{win}$ that achieves the optimal $\operatorname{aVal}(v_{0})$.
@@
In many applications, when a winning strategy does not exist, it is desirable for the Sys player to adopt a strategy that ensures the best possible outcome. For game $\mathcal{G}$ in Fig. 1(b), $\sigma_{win}$ does not exist but it is better for the Sys player to keep trying, i.e., move to $v_{1}$ or $v_{3}$ rather than giving up. To capture this intuition, we turn to the classical notion of dominance.
@@
Definition 5 (Dominance [Brenguier et al., 2016; Berwanger, 2007]).
Given two Sys strategies $\sigma,\sigma^{\prime}\in\Sigma$ and initial state $v\in V$, we say
@@
Brandenburger et al. [2008] rationalize a player playing admissibly to be doing their best. Thus, the first problem we consider is the synthesis of admissible strategies.
Problem 1.
Given a 2-player quantitative reachability game $\mathcal{G}$ and energy budget $\mathcal{B}\in\mathbb{N^{+}}$, synthesize the set of all admissible strategies $\Sigma_{adm}$ such that, for every $\sigma\in\Sigma_{adm}$, there exists $\tau\in\mathrm{T}$ under which $\operatorname{Val}(P^{v_{0}}(\sigma,\tau))\leq\mathcal{B}$.
Intuitively, $\sigma\in\Sigma_{adm}$ allows the Sys player to do its best without any assumptions about the Env player. In qualitative settings, winning strategies are always admissible [Brenguier et al., 2015, Lemma 8]. In quantitative settings, however, winning strategies are not necessarily admissible (e.g., $\sigma_{1}$ in Fig. 2). Thus, to compute enforceable admissible strategies, we introduce the notion of an admissible winning strategy.
@@
Strategy $\sigma$ is called admissible winning for Sys player iff it is admissible and $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$ if $\operatorname{aVal}(h)<\infty$ then $\operatorname{aVal}(h,\sigma)<\infty$.
For the game in Fig. 2, $\sigma_{3}$ is admissible but not enforcing as it cannot ensure always reaching the goal state. $\sigma_{1}$ is winning but not admissible. $\sigma_{2}$ is admissible winning, which is more desirable than the other two. In this work, we also consider the synthesis problem of such strategies.
Problem 2.
Given a 2-player quantitative reachability game $\mathcal{G}$ and energy budget $\mathcal{B}\in\mathbb{N}^{+}$, synthesize the set of all admissible winning strategies $\Sigma^{win}_{adm}$ such that, for every $\sigma\in\Sigma^{win}_{adm}$ and $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$, if $\operatorname{aVal}(h)<\infty$ then $\operatorname{aVal}(h,\sigma)\leq\mathcal{B}$.
In Section 3, we show how to solve Problem 1 by providing necessary and sufficient conditions for a strategy to be admissible. In Section 4, we identify the class of admissible strategies that are admissible winning and give an algorithm to solve Problem 2. Due to space constraints, the proofs of all theoretical claims are provided in Appendix.